# Generating quantum non-local entanglement with mechanical rotations

### v2

@@ I Introduction
Figure 1: (a) Photonic setting for the rotation-controlled generation of quantum non-local states of polarization. Two photons are initially prepared in the separable state $|HV\rangle$ and injected into the setup. The circulators, denoted by C, first send the photons into a Sagnac loop, and then redirect them towards two individual detection stages (only the directional paths, $1$ to $2$ and $2$ to $3$, are allowed by the circulators). As the photons entering the Sagnac loop through Beam-splitter BS are initially prepared in orthogonal polarization states, they do not interact at any point via electromagnetic couplings. We measure the polarization of the photons with a standard Bell detection scheme. (b) Theoretical prediction of the violation of the Bell-CHSH inequality. With Post-selection refers to the case where one photon is measured using the setup at the top (managed by Alice) and one photon is measured with the setup on the right (managed by Bob). Without Post-selection refers to the case where all the photons are considered in the analysis, which includes the detection of one photon by Alice and one of photon by Bob as well as of both photons by either Alice or Bob. For concreteness, we have set the photon wavelength to $\lambda=1\,\text{$\mu$m}$ ($\omega=2\pi c/\lambda$) and the interferometric area to $\sim 7.8\,\text{m}^{2}$ (e.g., $10$ loops of fiber with radius $r=0.5\,\text{m}$ with a total length of $\sim 31.5\,\text{m}$). We predict that a violation ($|{S}|>2$) occurs periodically with the rotation frequency as stated by Eq. (13). The maximum violation is first achieved for $\Omega_{\text{Bell}}\sim 0.4\,\text{Hz}$, and the condition $|S|>2$ is achieved in the interval $\Omega\in 0.4\pm 0.1$ Hz (see Appendix A for more details).
@@
$|\psi_{\text{i}}\rangle\rightarrow\ket{\psi_{1}}=\frac{1}{2}(\hat{a}_{H}^{\dagger}+i\hat{b}_{H}^{\dagger})(i\hat{a}_{V}^{\dagger}+\hat{b}_{V}^{\dagger})|0\rangle,$ (2)
@@
|$\ket{\psi_{1}}{\to}\ket{\psi_{2}}{=}\frac{1}{2}\left[e^{i\frac{\phi}{2}}\hat{a}_{H}^{\dagger}{+}ie^{-i\frac{\phi}{2}}\hat{b}_{H}^{\dagger}\right]\left[ie^{i\frac{\phi}{2}}\hat{a}_{V}^{\dagger}{+}e^{-i\frac{\phi}{2}}\hat{b}_{V}^{\dagger}\right]\!|0\rangle,$ (3)
@@
×$\times\left.\left(ie^{i\phi/2}(-i\hat{a}_{V}^{\dagger}+\hat{b}_{V}^{\dagger})+e^{-i\phi/2}(\hat{a}_{V}^{\dagger}-i\hat{b}_{V}^{\dagger}\right)\right]|0\rangle.$
@@ III Violation of the Bell-CHSH inequality
$|\psi\rangle$ is the state of the system, and $\hat{A}_{\bm{v}}$ and $\hat{B}_{\bm{u}}$ correspond to the measurements performed by Alice and Bob, respectively ($\bm{v}=\bm{a}$,$\bm{a}^{\prime}$ and $\bm{u}=\bm{b}$,$\bm{b}^{\prime}$ denote unit vectors parametrizing the measurement operators). The four terms appearing in Eq. (9) are assocated to four different joint measurements performed by Alice and Bob. Local realism enforces the Bell-CHSH inequality $|S|\leq 2$. However, as we will see, suitable choices of the angular frequency of rotation, encoded in the phase $\phi$ in Eq. (6), allows to violate such constraint. We show below that the Bell-CHSH inequality is violated up to the value $|S|=1+\sqrt{2}$ by considering the total state in Eq. (6), establishing the phenomena of rotationally induced quantum non-locality (Sec. III.1). Furthermore, we show that Tsirelson’s bound $|S|=2\sqrt{2}$ can be achieved by post-selecting coincidences at the detectors (Sec. III.2).
@@ III.1 Non-locality without post-selection
In particular, by setting $\phi=\pi/2+k\pi\leavevmode\nobreak\ (k\in\mathcal{Z})$, the state in Eq. (6) reduces to $|\psi_{\text{tot}}\rangle=\frac{1}{\sqrt{2}}(\psi_{\text{\text{\text{fav}}}}\rangle+|\psi_{\text{\text{\text{unfav}}}}\rangle)$, i.e.,
@@ III.2 Tsirelson’s bound with postselection
In particular, as stated in Eq. (15), by setting $\phi=\pi/2+k\pi\leavevmode\nobreak\ (k\in\mathcal{Z})$, the state in Eq. (14) reduces to the Bell state $|\Psi^{-}\rangle$ and we achieve the notorious Tsirelson’s bound $|S|=2\sqrt{2}$ [41]. In Fig. 1(b), we plot the Bell-CHSH function defined in Eq. (16) for the same set of values as for the case without postselection to ease the comparison.
@@ Appendix C Effect of Noise on Bell-CHSH Test
where $\hat{\rho}_{AB}=\ket{\psi_{3}}\bra{\psi_{3}}$ |, with |$\ket{\psi_{3}}$ the interferometer output state give by Eq. (6), while $\hat{\rho}_{A}=\text{tr}_{B}\left(\ket{\psi_{3}}\bra{\psi_{3}}\right)$ and $\hat{\rho}_{B}=\text{tr}_{A}\left(\ket{\psi_{3}}\bra{\psi_{3}}\right)$ are the reduced states of Alice’s and Bob’s subsystem. We use $\hat{\rho}$ to compute the Bell-CHSH function $S$ from Eq. (9) for two cases: without post-selection (Sec. III.1) and with post-selection (Sec. III.2).

### v3

@@ I Introduction
Figure 1: (a) Photonic setting for the rotation-controlled generation of quantum non-local states of polarization. Two photons are initially prepared in the separable state $|HV\rangle$ and injected into the setup. The circulators, denoted by C, first send the photons into a Sagnac loop, and then redirect them towards two individual detection stages (only the directional paths, 11 to $2$ and 22 to $3$, are allowed by the circulators). As the photons entering the Sagnac loop through Beam-splitter BS are initially prepared in orthogonal polarization states, they do not interact at any point via electromagnetic couplings. We measure the polarization of the photons with a standard Bell detection scheme. (b) Theoretical prediction of the violation of the Bell-CHSH inequality. With Post-selection refers to the case where one photon is measured using the setup at the top (managed by Alice) and one photon is measured with the setup on the right (managed by Bob). Without Post-selection refers to the case where all the photons are considered in the analysis, which includes the detection of one photon by Alice and one of photon by Bob as well as of both photons by either Alice or Bob. For concreteness, we have set the photon wavelength to $\lambda=1\,\text{$\mu$m}$ ($\omega=2\pi c/\lambda$) and the interferometric area to $\sim 7.8\,\text{m}^{2}$ (e.g., 1010 loops of fiber with radius $r=0.5\,\text{m}$ with a total length of $\sim 31.5\,\text{m}$). We predict that a violation (|S|>2|{S}|>2) occurs periodically with the rotation frequency as stated by Eq. (13). The maximum violation is first achieved for $\Omega_{\text{Bell}}\sim 0.4\,\text{Hz}$, and the condition $|S|>2$ is achieved in the interval $\Omega\in 0.4\pm 0.1$ Hz (see Appendix A for more details).
@@
[…]
@@
[…]
@@
$\times\left.\left(ie^{i\phi/2}(-i\hat{a}_{V}^{\dagger}+\hat{b}_{V}^{\dagger})+e^{-i\phi/2}(\hat{a}_{V}^{\dagger}-i\hat{b}_{V}^{\dagger}\right)\right]|0\rangle.$
@@ III Violation of the Bell-CHSH inequality
$|\psi\rangle$ is the state of the system, and $\hat{A}_{\bm{v}}$ and $\hat{B}_{\bm{u}}$ correspond to the measurements performed by Alice and Bob, respectively ($\bm{v}=\bm{a}$$\bm{a}^{\prime}$ and $\bm{u}=\bm{b}$$\bm{b}^{\prime}$ denote unit vectors parametrizing the measurement operators). The four terms appearing in Eq. (9) are assocated to four different joint measurements performed by Alice and Bob. Local realism enforces the Bell-CHSH inequality $|S|\leq 2$. However, as we will see, suitable choices of the angular frequency of rotation, encoded in the phase $\phi$ in Eq. (6), allows to violate such constraint. We show below that the Bell-CHSH inequality is violated up to the value $|S|=1+\sqrt{2}$ by considering the total state in Eq. (6), establishing the phenomena of rotationally induced quantum non-locality (Sec. III.1). Furthermore, we show that Tsirelson’s bound $|S|=2\sqrt{2}$ can be achieved by post-selecting coincidences at the detectors (Sec. III.2).
@@ III.1 Non-locality without post-selection
In particular, by setting $\phi=\pi/2+k\pi\penalty 10000\ (k\in\mathcal{Z})$, the state in Eq. (6) reduces to $|\psi_{\text{tot}}\rangle=\frac{1}{\sqrt{2}}(\psi_{\text{\text{\text{fav}}}}\rangle+|\psi_{\text{\text{\text{unfav}}}}\rangle)$, i.e.,
@@ III.2 Tsirelson’s bound with postselection
In particular, as stated in Eq. (15), by setting $\phi=\pi/2+k\pi\penalty 10000\ (k\in\mathcal{Z})$, the state in Eq. (14) reduces to the Bell state $|\Psi^{-}\rangle$ and we achieve the notorious Tsirelson’s bound $|S|=2\sqrt{2}$ [41]. In Fig. 1(b), we plot the Bell-CHSH function defined in Eq. (16) for the same set of values as for the case without postselection to ease the comparison.
@@ Appendix C Effect of Noise on Bell-CHSH Test
[…]